Solve by the substitution method
5x+3y=2
x=22-3y

Answer by Jk22(389)

5x+3y=2
x=22-3y | subs x in preceding equation :

5*(22-3y)+3y=2 => 110-15y+3y=2 => 12y=108 => 6y=54 => y=9, x=22-3*9=22-27=-5

Verification :
5*(-5)+3*9=-25+27=2
22-3*9=22-27=-(27-22)=-5=x